What is (a) the empirical formula and (b) the molecular formula of a compound (molecular mass=60.10 g/mol) containing carbon and hydrogen? A lab technician combusts the 0.7148 g sample and it produced 1.048x10^3 mg CO2 and 8.573x10^-1 g pf hydrogen.
Answer #1

mass of CO2 = 1.048 g

moles of CO2 = 1.048 / 44

                      = 0.0238

moles of C = 0.0238

moles of H2 = 0.8573 / 2 = 0.4286

mole of H = 0.4286

moles ratio = 0.0238 : 0.4286

                  = 1 : 18

a) empirical formula   = CH18

empirical formula = 12 + 18 x 1 = 30

n = 60 / 30 = 2

molecular formula   = 2 x C H18 = C2H36

b) molecular formula = C2H36